# The FRW Universe as a van der Waals-like Thermodynamic Heat Engine

### v1

@@ Abstract
It is well known that the Friedmann-Robertson-Walker (FRW) universe is a dynamical spacetime, and it has thermodynamics embodied on the apparent horizon. Notably, it also possesses a van der Waals-like equation of state, enabling us to consider thermodynamic cycles and explore the FRW universe’s potential as a heat engine. In this paper, we investigate two types of cycles–the Carnot cycle and the rectangular cycle–based on the phase diagram derived from the equation of state, to study the heat engine characteristics of the FRW universe. Furthermore, we calculate the work done and assess the corresponding efficiencies, illustrating the efficiency diagram for the FRW universe’s heat engine. We observe that the efficiency of the rectangular cycle consistently remains below unity and never exceeds the Carnot efficiency–the thermodynamic upper limit. This finding aligns with traditional thermodynamic principles governing heat engines.
@@ I Introduction
Black hole thermodynamics is a fruitful area in modern theoretical physics, which has inspired several cutting edge studies. Black hole phase transition and critical phenomenon takes a pivotal status in black hole thermodynamics. In explorations of critical phenomenon and phase transition of black holes, the thermodynamic pressure plays a key role. In recent years, the cosmological constant $\Lambda$ is treated as the thermodynamic variable analogous to the pressure $P_{\Lambda}=-{\Lambda}/{8\pi}$ Kastor:2009wy ; Dolan:2010ha and its conjugate quantity is the thermodynamic volume, which yields the presence of a pressure-volume term in the first law of thermodynamics. In this framework, the idea of pressure and volume as well as the thermodynamic equation of state in the extended phase space of black hole thermodynamics in Anti-de Sitter (AdS) spacetime has drawn a lot of attention Kubiznak:2012wp ; Kubiznak:2016qmn ; Wei:2012ui ; Gunasekaran:2012dq ; Hendi:2012um ; Cai:2013qga ; Altamirano:2014tva ; Dehghani:2014caa ; Xu:2015rfa ; Cheng:2016bpx ; Hu:2018qsy ; Estrada:2019cig ; Li:2020xkh ; Abdusattar:2023xxs ; Abdusattar:2024zzi .
@@ III.2 The Rectangular Cycle
$C_{P}$ $=$ $\left(\frac{\partial{\cal H}}{\partial T}\right)_{P}=T\left(\frac{\partial S}{\partial T}\right)_{P}$ (23)
$=$ $\frac{2{\pi}R_{A}^{2}(-1+8{\pi}R_{A}^{2}P)}{1+8{\pi}R_{A}^{2}P}=\frac{4{\pi}^{2}R_{A}^{3}T}{1+2{\pi}TR_{A}}\,$, (24)
@@ V Acknowledgment
This work is supported by National Natural Science Foundation of China (NSFC) under grant No. 12465012, Kashi University high-level talent research start-up fund project under grant No. 022024002, and Tianchi Talented Young Doctors Program of Xinjiang Uyghur Autonomous Region.

### v2

# footnotetext: $^{\text{\Letter}}$ Corresponding Author

@@ Abstract
It is well known that the Friedmann-Robertson-Walker (FRW) universe is a dynamical spacetime, and it has thermodynamics embodied on the apparent horizon. Notably, it also possesses a van der Waals-like equation of state, enabling us to consider thermodynamic cycles and explore the potential of the FRW universe as a heat engine. In this paper, we investigate two types of cycle–the Carnot cycle and the rectangular cycle–based on the phase diagram derived from the equation of state, to study the heat engine characteristics of the FRW universe. Furthermore, we calculate the work done and assess the corresponding efficiencies, illustrating the efficiency diagram for the FRW universe’s heat engine. We observe that the efficiency of the rectangular cycle consistently remains below unity and never exceeds the Carnot efficiency–the thermodynamic upper limit. This finding is in alignment with the traditional thermodynamic principles that govern heat engines.
@@ I Introduction
Black hole thermodynamics is a fruitful area in modern theoretical physics, which has inspired several cutting edge studies. Black hole phase transition and critical phenomenon takes a pivotal status in black hole thermodynamics. In explorations of critical phenomenon and phase transition of black holes, the thermodynamic pressure plays a key role. In recent years, the cosmological constant $\Lambda$ is treated as the thermodynamic variable analogous to the pressure $P_{\Lambda}=-{\Lambda}/{8\pi}$ Kastor:2009wy ; Dolan:2010ha and its conjugate quantity is the thermodynamic volume, which yields the presence of a pressure-volume term in the first law of thermodynamics. In this framework, the idea of pressure and volume as well as the thermodynamic equation of state in the extended phase space of black hole thermodynamics in Anti-de Sitter (AdS) spacetime has drawn a lot of attention Kubiznak:2012wp ; Kubiznak:2016qmn ; Wei:2012ui ; Gunasekaran:2012dq ; Hendi:2012um ; Cai:2013qga ; Altamirano:2014tva ; Dehghani:2014caa ; Xu:2015rfa ; Cheng:2016bpx ; Hu:2018qsy ; Estrada:2019cig ; Li:2020xkh ; Abdusattar:2023xxs ; Abdusattar:2024zzi ; Abdusattar:2025rdp .
@@
At this point, it is important to emphasize that the thermodynamic interpretation adopted throughout this work should be understood as an effective description. In a globally dynamical spacetime such as the FRW universe, notions like temperature, pressure, work, and thermodynamic cycles are expected to remain physically meaningful only within regimes where the apparent horizon evolves sufficiently slowly and quasi-equilibrium conditions can be assumed. Outside such regimes, thermodynamic quantities may still be formally defined, but their physical interpretation becomes conditional rather than global. The heat-engine picture explored in this paper should therefore be interpreted within this domain of applicability.
@@ III Thermodynamic Heat Engine Cycle Analysis in the FRW Universe
It is worth clarifying that the thermodynamic cycles considered here should be interpreted as effective paths in the space of thermodynamic variables, rather than as literal cyclic evolutions of the cosmological spacetime. The heat-engine description relies on the assumption that, along these paths, thermodynamic quantities such as pressure, volume, and temperature remain operationally well defined. In a fully dynamical FRW background, this interpretation is therefore expected to apply within regimes where the evolution is sufficiently slow for a quasi-equilibrium description to hold.
@@ III.2 The Rectangular Cycle
$C_{P}$ $=$ $\left(\frac{\partial{\cal H}}{\partial T}\right)_{P}=T\left(\frac{\partial S}{\partial T}\right)_{P}$
$=$ $\frac{2{\pi}R_{A}^{2}(-1+8{\pi}R_{A}^{2}P)}{1+8{\pi}R_{A}^{2}P}=\frac{4{\pi}^{2}R_{A}^{3}T}{1+2{\pi}TR_{A}}\,$,
@@ IV Conclusions and Discussion
Finally, it is worth emphasizing that the thermodynamic and heat-engine interpretation presented in this work should be understood as an effective description, whose physical applicability depends on the dynamical regime of the FRW universe. The efficiencies, cycles, and bounds discussed here are expected to be meaningful within phases where a quasi-equilibrium thermodynamic description associated with the apparent horizon remains operational. Explicitly acknowledging this domain of validity helps frame the results in a consistent physical context and clarifies the scope within which the heat-engine picture can be reliably applied.
Beyond the quasi-equilibrium regime implicitly assumed throughout this analysis, departures from the expected thermodynamic relations should be interpreted not as violations of thermodynamic principles, but as indications of the breakdown of the effective thermodynamic description itself. In such situations, quantities such as temperature, entropy, pressure, volume, or even thermodynamic cycles and efficiencies may remain formally definable while progressively losing a clear operational meaning associated with physical heat exchange and work. From this perspective, the limits of applicability of the heat-engine picture correspond to physically meaningful transitions in the dynamical behavior of the FRW universe, rather than inconsistencies of the framework. This highlights the importance of viewing the present construction as an effective description with a well-defined domain of validity.
@@ V Acknowledgment
The authors thank Profs. Y. P. Hu, Hongsheng Zhang, and Dr. S. B. Kong for their insightful discussions. This work is supported by the National Natural Science Foundation of China (NSFC) under grant No. 12465012, the Kashi University high-level talent research start-up fund project under grant No. 022024002, and the Tianchi Talented Young Doctors Program of the Xinjiang Uyghur Autonomous Region.